perimeter of a rectangle is 78 m. The length is 9 m more than twice the width. Find the dimentions of the lenghth and width.
Answer by checkley75(3666)

2L+2W=PERIMETER IS THE FORMULA.
L=2W+9 THIS IS GIVEN
2(2W+9)+2W=78
4W+18+2W=78
6W=78-18
6W=60
W=60/6
W=10 m. ans. for the width.
L=2*10+9
l=20+9
L=29 m. ans. for the length.
Proof:
2*29+2*10=78
58+20=78
78=78